OpenStudy (anonymous):

E=mc^2, where E is energy (in joules), m is mass (in kg) and c is the speed of light (c=3.0x10^8 m/s). A penny has a mass of 2.35g. How much energy would be produced if it was converted into energy? Show the steps and state the final answer in kWh.

OpenStudy (anonymous):

Just plug the mass (in kg) into the formula. You will get energy in joules. After you get energy in joules, convert it into kWh. (1 J = 1 Ws)

OpenStudy (anonymous):

1 kWh = 10^3*3600 Ws =3.6*10^6 Ws

OpenStudy (anonymous):

oh, okay thanks!